The following linear regression equation is for the age (year) and price ($) of used Ford Escorts. P r i c e space equals space 8 comma 857 space minus 1 comma 114 A g e Predict the price of a 3 year old Ford Escort.
Answer #1

The given linear regression equation is for the age (year) and price ($) of used Ford Escorts.

Price = 8,857 - 1,114 Age

Plug Age = 3 in the above equation , we get

Price = 8,857 - ( 1,114*3 ) = 8,857 - 3,342 = 5,515

So we predict the price of a 3 year old Ford Escort is as $5,515